# Lambda-ABF-OPES: Faster Convergence with High Accuracy in Alchemical Free Energy Calculations

### v1

@@ Abstract
1 Introduction
Accurately predicting the binding affinity between a small molecule (ligand) and its target (protein or RNA/DNA) is a fundamental tool in modern computational drug discovery, that can be used both at the hit discovery and at the lead optimization stages 1. Indeed, it minimizes the cost of such project by reducing the need to rely on experimental assays.
@@ 1 Introduction
To overcome this limitation, we combine Lambda-ABF with the exploratory version of the OPES method, known as OPES-Explore 27. Hereafter, we will simply refer to it as OPES. This method is CV-based and aims at sampling a target probability distribution associated to lower barriers. It can be seen as an evolution of (well-tempered) MetaDynamics with an emphasis on exploration and is user-friendly as it requires only a few physically motivated parameters to be set. By applying a bias to the lambda CV within the Lambda-ABF-OPES method, we effectively integrate the strengths of both ABF and OPES.
Similarly to the (wt)meta-eABF method 29, the Lambda-ABF-OPES approach benefits from the combination of ABF and OPES. While ABF reduces the free energy barriers, OPES fills the energy valleys by incorporating a history-dependent potential term. To our knowledge, this is the first instance of combining ABF with the OPES method in an alchemical framework. This approach improves convergence speed, enhancing it by up to 9 times compared to the original Lambda-ABF method.
We applied this hybrid approach, utilizing the AMOEBA polarizable force field and DBC restraints, to calculate the absolute binding free energy of 11 diverse small drug-like molecule inhibitors binding to bromodomains (BRD4), a well-studied macromolecular system 39, 40, 41, 42. Our results show that while maintaining high accuracy, with a mean absolute error of 0.90 kcal/mol, we achieved a significant improvement in convergence speed. This represents a notable advancement, particularly when using the more computationally demanding AMOEBA polarizable force field.
@@
2 Methods
2.1 MD Simulations
We performed all simulations using the molecular dynamics code Tinker-HP Version 1.2 43, 44, which includes the Colvars library 37 for Lambda-ABF-OPES simulations. For the protein, water, and ion parameters, the polarizable AMOEBA force field 9, 10, 11, 12 was applied. Ligand parameterization was performed using the Poltype package 13. The systems, containing between 31,000 and 42,000 atoms on average (varying with ligand size), were neutralized with 0.15 M physiological NaCl concentration. A cubic water box with dimensions of 70–80 Å was used, ensuring a minimum distance of 12 Å between the ligand and the box edge, with its size adjusted according to the ligand.
The systems were first minimized using the Tinker-HP minimize program, followed by a two-step heating process from 200 K to 298 K under the NVT ensemble. In the first step, restraints were applied to all protein-ligand and X-ray water molecule atoms. In the second step, restraints were applied to heavy backbone atoms, using a force constant of 10 in both steps. Each stage included 4 ns of MD simulations, conducted using the RESPA integrator with a time step of 2.0 fs.
In the third step, the three-level multiple timestep BAOAB-RESPA1 integrator 45 with an outer timestep of 10 fs (and an intermediate of 3.33 fs and a shorter timestep of 1 fs) was employed under the NVT ensemble for 5 ns, with restraints applied only to C atoms using a force constant of 1 . This was followed by equilibration in the NPT ensemble at 1 atm for 5 ns, maintaining the same restraints as the previous step. Finally, for each ligand, a production run of at least 100 ns was performed under the NPT ensemble using the BAOAB-RESPA1 Langevin integrator with a 10 fs time step. During the production phase, a restraint with a force constant of 1 was applied to the C atoms of the protein’s floppy tail to prevent large conformational changes.
Non-Langevin temperature control was ensured by using the Bussi thermostat 46 and pressure with the Berendsen barostat 47. Van der Waals interactions employed a 12 Å cutoff, while electrostatic interactions were treated using the Particle Mesh Ewald (PME) method 48 with a real-space cutoff of 7 Å. Induced dipoles were calculated with a Preconditioned Conjugate Gradient (PCG) solver, with a convergence tolerance of Debye 49.
2.2 DBC Restraint
@@
To select the LA, we monitored the root mean square fluctuation (RMSF) of the ligand’s heavy atoms over at least 100 ns of standard MD simulations. Atoms with an RMSF of less than 0.6 Å were selected for tight binders (Ligands 1-9), while those with an RMSF between 0.7 and 0.8 Å were chosen for weak binders (Ligands 10-11). For the RB selection, we used the C atoms of the protein within 6 Å of the ligand, with an RMSF of approximately 0.6 Å. See Fig. S1 in the SI for the definition of DBC for the protein and each ligand. During plain MD of each ligand, the DBC was monitored using the Colvars library, and the DBC value within the 95% interval of the distribution was selected as the DBC cutoff for the restraint in Lambda-ABF-OPES. This selection ensures that during the alchemical simulation, there is no biased artifact from the DBC restraint. The DBC cutoff for each ligand is listed in Table S3 SI. A flat-bottomed harmonic restraint was then applied to the DBC with a force constant of 100 above the DBC cutoff.
2.3 Lambda-ABF-OPES for Calculation of Absolute Binding Free Energies
@@
Details of the Lambda-ABF and OPES-Explore methods can be found in the original publications 34, 26, 27. Here we provide a brief overview of their underlying theories.
@@
The Lambda-ABF algorithm 21, 22, 34 adaptively computes the derivative of the free energy associated with the parameter using the thermodynamic integration (TI) formula 50. This estimate is applied as a force directly to the simulated system, guiding the dynamics. As a result, the sampling of converges toward a uniform distribution, facilitating the crossing of free-energy barriers.
@@
OPES-Explore 27, a collective-variable (CV)-based enhanced sampling technique, represents the latest advancement in the MetaDynamics family. It broadens the sampling of a system toward a target probability distribution, known as the well-tempered distribution. This approach employs Gaussian kernels to adaptively construct a bias potential, guiding the system’s exploration while preserving thermodynamic relevance. Critical parameters, such as the Gaussian kernel width and bias factor, are pivotal in defining the sampling’s scope and stability. Additionally, the barrier parameter () ensures efficient transitions between basins while preventing access to irrelevant high-energy states.
@@
In this study, the ABFE is calculated by continuously “alchemically” decoupling the ligand from its environment, both in complex with the protein and separately in bulk solvent. The standard free energy of binding is then determined using a thermodynamic cycle 51, which requires sampling of the alchemical Hamiltonians. We employed the Lambda-ABF approach as implemented in Tinker-HP/Colvars, in combination with OPES-Explore provided by the Colvars library 37 version 2024-11-18. This integration provides user-friendly convergence estimation without requiring post-processing and allows seamless compatibility with other CV-based methods.
@@
All simulations were performed at T = 300 K and P = 1 Atm using the BAOAB-RESPA integrator 45 with a 3 fs time step under the NPT ensemble, employing four walkers 52 to separately decouple the VDW and ELE interactions. More precisely, the polarizabilities and permanent multipoles of the ligand are scaled down to 0 in the electrostatic legs (ELE), and the van der Waals interactions between the atoms of the ligand and all the other ones are scaled down to 0 (leveraging softcore interactions 53) in the van der Waals legs (VDW). In the complex phase, 30 ns were simulated for the VDW leg and 5 ns for the ELE leg. In the solvent phase, both the VDW and ELE legs were simulated for 5 ns each.
@@
As mentioned earlier, for OPES-Explore, the barrier parameter is a critical setting in the simulations and must be sufficient to ensure that the method accurately captures the energy landscape and transition state barriers. We conducted tests using OPES-Explore alone with different barriers. The final values varied between runs (see SI), indicating that OPES-Explore alone struggles to compute the binding free energies of interest. However, when combining OPES-Explore with Lambda-ABF, the system is primarily driven by Lambda-ABF, which facilitates the crossing of energy barriers, and also benefits from the rapid convergence of the TI estimator. In this case, OPES-Explore mainly serves to push the system out of local minima. Therefore, the barrier parameter does not need to be equal to the free energy differences associated to the simulation. Our tests showed that setting the bias threshold of OPES-Explore to a minimum of 5 kcal/mol for both the ELE and VDW legs in the complex and solvent phases effectively accelerated the convergence across all ligands, regardless of their respective energy barriers. For ligand 11, a threshold of 2 kcal/mol was sufficient for proper acceleration. Although a barrier of 5 kcal/mol also works well, since this ligand is a weak binder, the lower threshold of 2 kcal/mol ensures better stability in convergence. An adaptive sigma is used to determine the Gaussian kernel width, and the frequency for kernel deposition was set to 300 steps, which corresponds to 9 ps with a time step of 3 fs.
For Ligand 11, during the 180 ns plain MD simulations, we observed two equally plausible binding modes in which the trifluorotoluene moiety was flipped by , consistent with the result of Ref. 41. Consequently, ABFE calculation was performed for each binding mode, resulting in a total of two binding free energy calculations. The results from multiple binding modes were combined into a single binding free energy value as described in Ref. 54.
@@
The free energy cost associated with the release of the DBC restraint is computed in the gas phase through TI by progressively releasing it to a compatible harmonic distance restraint which can then be computed analytically 34, 51, 36.
@@
In these simulations, we observed that the rotamers of Asn140 in the Apo state may adopt a conformation favorable in the Apo state but not in the Holo state (see Results and Discussion and SI for more details). This could lead to sampling the Holo state with an incorrect orientation of Asn140, potentially introducing artifacts. The restriction (through a restraint) to one of these rotamers as an Apo endpoint must be accounted for. Therefore, a positional restraint was applied to the heavy atoms of the Asn140 residue and its neighboring atoms (all heavy atoms of the backbone and those within 6 Å of the ligand with a mild force constant of 2 ). This restraint was necessary to ensure that Asn140 and its neighboring atoms were properly sampled in both the Apo and Holo states, given the continuous switching of lambda between 0 (Apo) and 1 (Holo) in the Lambda-ABF-OPES simulations. Since both rotamers of Asn140 are equally favorable in the Apo state (see Fig. 2), an RTln(2) correction was added to the final computed . In addition, to avoid artifacts during alchemical decoupling as described in SI, the C atoms of BRD4 were restrained to their relaxed configuration.
2.4 Analytical Lambda Derivatives for Variational Many-Body Potentials
@@
As all TI-based technique, Lambda-ABF requires the computation of potential derivatives with respect to the alchemical parameter which are not trivial to compute for many-body interactions such as polarization as present in the AMOEBA force field. In our previous work34 we used a simple interpolation of polarization between the end states which gives immediately the associated derivatives as the difference of these, but this is associated with an increase of the computational cost because of the need to solve 2 polarization equations at each timestep. Recently, analytical derivatives associated to a simple scaling of the polarizabilities and multipoles have been introduced55. Here we resorted to a more general formulation relying on the variational formulation of the many-body term and the Hellman-Feynman theorem: and
@@
because of the minimum conditions on the induced dipoles. Note that this formulation still holds for other many-body terms with similar variational formulation as is the case for fluctuating charges56, continuum solvation models57 or QM-MM58, 59. Further development of the gradients used in this work using Particle Mesh Ewald is given in Technical Appendix.
3 Results and Discussion
@@
In the following section, we will first provide a brief overview of the general binding modes of the 11 ligands in the BRD4 system. We will then focus on the orientation of the Asn140 residue, a key residue that directly interacts with the ligands, in both the Apo and Holo states, highlighting any significant differences. Finally, we will examine the binding affinities and correlate these findings with the experimental values.
@@
As depicted in Fig. 1(c), the 11 ligands are large, flexible, drug-like molecules, some of which are charged (Ligand 1 and 4). These characteristics make them an ideal set for evaluating the performance of the new method. All ligands target a common binding site. Asparagine (Asn140), located within the BC-loop binding pocket (Fig. 1(a)), is the most critical residue, which interacts directly with the ligands. The side chain of Asn140 consists of an amide group () attached to a C, which is itself connected to the C of the backbone via rotatable bonds.
@@
In addition to Asn140, four conserved (polarizable) water molecules 60 in the binding site of BRD4 also play a crucial role in stabilizing the ligands within the binding pocket, thereby enhancing binding affinities. As illustrated in Fig. 1(a), one of these water molecules forms a bridge between the Tyr97 residue and the ligand (except for Ligand 4). To preserve this stabilization, the X-ray water molecules were retained in the binding pocket during system preparation, as they contribute to a hydrogen-bond network involving the three other conserved water molecules and the protein.
Using the equilibrated structures, we carried out absolute binding free energy calculations employing the novel Lambda-ABF-OPES method. The correlation plot between experimental values and calculated results is shown in Fig. 3 and reported in Table 1.
We observe a strong correlation between the experimental and calculated results, with a Pearson’s correlation coefficient (Pearson’s ) of 0.81. The analysis yielded a root mean square error (RMSE) of 1.1 kcal/mol and a mean absolute error (MAE) of 0.9 kcal/mol. These results are consistent with experimental data and align well with those reported in Refs. 41, 40. The key advantage of this method lies in its user-friendliness, reduced computational resource requirements, and rapid convergence. Compared to Lambda-ABF, the traditional fixed Lambda approach 41, and the newly developed CV-based approach 40, this method offers significantly improved computational efficiency while maintaining high accuracy. These attributes make it particularly well-suited for large-scale or high-throughput applications.
@@
Figure 4 illustrates an example of a convergence plot over time for the ELE and VDW legs of the Ligand 8 complex phase using Lambda-ABF and Lambda-ABF-OPES method in a single replica. In this case, the acceleration in convergence time is a factor of 9 for ELE and 4 for VDW. When considering different replicas, the convergence speedup for ELE ranges from 5 to 9 times, while for VDW, it ranges from 3 to 5 times.
@@ 4 Conclusions
In this study, we introduced a new hybrid approach that combines Lambda-ABF with the exploratory version of the OPES method. This novel integration leverages the complementary strengths of ABF and OPES to overcome critical limitations in alchemical free energy calculations, including inefficient exploration of configurational space and kinetic trapping in energy landscapes. By applying biases to the Lambda collective variable (CV) and incorporating the AMOEBA polarizable force field alongside DBC restraints, our method achieves unprecedented levels of sampling efficiency, up to nine times faster than the original Lambda-ABF technique.
Our application of this hybrid method to a diverse set of 11 drug-like molecules targeting BRD4 bromodomains yielded close alignment between our computational results and experimental data with a mean absolute error of 0.9 kcal/mol. Importantly, this was achieved while using the highly accurate AMOEBA polarizable force field, demonstrating the feasibility of this approach for real-world drug discovery applications. This approach can be naturally extended to neural networks methodologies including Machine Learning Interatomic Potentials (MLIP) 68, 69 and Foundation models whose additional computational cost for free energy computations compared to FFs has limited to date their use in production.

### v2

@@ Abstract
Accurately predicting the binding affinity between a small molecule (ligand) and its target (protein or RNA/DNA) is a fundamental tool in modern computational drug discovery, that can be used both at the hit discovery and at the lead optimization stages 1. Indeed, it minimizes the cost of drug discovery by reducing the need to rely on experimental assays.
@@
To overcome this limitation, we combine Lambda-ABF with the exploratory version of the OPES method, known as OPES-Explore 27 (OPESe). This method is CV-based and aims at sampling a target probability distribution associated to lower barriers. It can be seen as an evolution of (well-tempered) MetaDynamics with an emphasis on exploration and is user-friendly as it requires only a few physically motivated parameters to be set. By applying a bias to the lambda CV within the Lambda-ABF-OPES method, we effectively integrate the strengths of both ABF and OPESe.
Similarly to the (wt)meta-eABF method 29, the Lambda-ABF-OPES approach benefits from the combination of ABF and OPESe. While ABF reduces the free energy barriers, OPESe fills the energy valleys by incorporating a history-dependent potential term. To our knowledge, this is the first instance of combining ABF with the OPESe method in an alchemical framework. This approach improves convergence speed, enhancing it by up to 9 times compared to the original Lambda-ABF method.
We applied this hybrid approach, utilizing the AMOEBA polarizable force field and DBC restraints, to calculate the absolute binding free energy of 11 diverse small drug-like molecule inhibitors binding to bromodomains (BRD4) as well as the Benzamidine-Trypsin complex, a well-studied macromolecular systems 39, 40, 41, 42, 43. Our results show that while maintaining high accuracy, with a mean absolute error of 0.90 kcal/mol for the BRD4 system, we achieved a significant improvement in convergence speed. For the Benzamidine-Trypsin complex, we also obtained very close agreement with the experimental value, with a notably low computational cost compared to, for example, the results of Ref.43. This represents a notable advancement, particularly when using the more computationally demanding AMOEBA polarizable force field.
@@
In the case of BRD4 complexes, to select the LA, we monitored the root mean square fluctuation (RMSF) of the ligand’s heavy atoms over at least 100 ns of standard MD simulations. Further details on simulation setup can be found in Supporting Information. Atoms with an RMSF of less than 0.6 Å were selected for tight binders (Ligands 1-9), while those with an RMSF between 0.7 and 0.8 Å were chosen for weak binders (Ligands 10-11). For the Benzamidine-Trypsin complex, due to the small size of the ligand, all heavy atoms of the ligand were selected for the DBC.
@@
For the RB selection of BRD4 and Trypsin, we used the C atoms of the protein within 6 Å of the ligand, with an RMSF below approximately 0.6 Å. See Fig. S1 in the SI for the definition of DBC for the protein and each ligand. During plain MD of each ligand, the DBC was monitored using the Colvars library, and the DBC value within the 95% interval of the distribution was selected as the DBC cutoff for the restraint in Lambda-ABF-OPES. This selection ensures that during the alchemical simulation, there is no biased artifact from the DBC restraint. The DBC cutoff for each ligand is listed in Table S1 SI. A flat-bottomed harmonic restraint was then applied to the DBC with a force constant of 100 above the DBC cutoff.
@@
Details of the Lambda-ABF and OPESe methods can be found in the original publications 34, 26, 27. Here we provide a brief overview of their underlying theories.
@@
The Lambda-ABF algorithm 21, 22, 34 adaptively computes the derivative of the free energy associated with the parameter using the thermodynamic integration (TI) formula 44. This estimate is applied as a force directly to the simulated system, guiding the dynamics. As a result, the sampling of converges toward a uniform distribution, facilitating the crossing of free-energy barriers.
@@
OPESe 27, a collective-variable (CV)-based enhanced sampling technique, represents the latest advancement in the MetaDynamics family. It broadens the sampling of a system toward a target probability distribution, known as the well-tempered distribution. This approach employs Gaussian kernels to adaptively construct a bias potential, guiding the system’s exploration while preserving thermodynamic relevance. Critical parameters, such as the Gaussian kernel width and bias factor, are pivotal in defining the sampling’s scope and stability. Additionally, the barrier parameter () ensures efficient transitions between basins while preventing access to irrelevant high-energy states.
@@
In this study, the ABFE is calculated by continuously “alchemically” decoupling the ligand from its environment, both in complex with the protein and separately in bulk solvent. The standard free energy of binding is then determined using a thermodynamic cycle 45, which requires sampling of the alchemical Hamiltonians. We employed the Lambda-ABF approach as implemented in Tinker-HP/Colvars, in combination with OPESe provided by the Colvars library 37 version 2024-11-18. This integration provides user-friendly convergence estimation without requiring post-processing and allows seamless compatibility with other CV-based methods.
@@
All simulations were performed at T = 300 K and P = 1 Atm using the BAOAB-RESPA integrator 46 with a 3 fs time step under the NPT ensemble, employing four walkers 47 to separately decouple the VDW and ELE interactions. More precisely, the polarizabilities and permanent multipoles of the ligand are scaled down to 0 in the electrostatic legs (ELE), and the van der Waals interactions between the atoms of the ligand and all the other ones are scaled down to 0 (leveraging softcore interactions 48) in the van der Waals legs (VDW). In the complex phase, 30 ns were simulated for the VDW leg and 5 ns for the ELE leg. In the solvent phase, both the VDW and ELE legs were simulated for 5 ns each.
@@
As mentioned earlier, for OPESe, the barrier parameter is a critical setting in the simulations and must be sufficient to ensure that the method accurately captures the energy landscape and transition state barriers. We conducted tests using OPESe alone with different barriers. The final values varied between runs (see SI Fig. S20), indicating that OPESe alone struggles to compute the binding free energies of interest. However, when combining OPESe with Lambda-ABF, the system is primarily driven by Lambda-ABF, which facilitates the crossing of energy barriers, and also benefits from the rapid convergence of the TI estimator. In this case, OPESe mainly serves to push the system out of local minima. Therefore, the barrier parameter does not need to be equal to the free energy differences associated to the simulation. Our tests showed that setting the bias threshold of OPESe to a maximum of 5 kcal/mol for both the ELE and VDW legs in the complex and solvent phases effectively accelerated the convergence across all ligands in the BRD4 and Trypsin systems, regardless of their respective energy barriers. For ligand 11 in complex with BRD4, a threshold of 2 kcal/mol was sufficient for proper acceleration. Although a barrier of 5 kcal/mol also works well, since this ligand is a weak binder, the lower threshold of 2 kcal/mol ensures better stability in convergence. An adaptive sigma is used to determine the Gaussian kernel width, and the frequency for kernel deposition was set to 300 steps, which corresponds to 9 ps with a time step of 3 fs.
@@
The free energy cost associated with the release of the DBC restraint is computed in the gas phase through TI by progressively releasing it to a compatible harmonic distance restraint which can then be computed analytically 34, 45, 36.
@@
In the BRD4 simulations, we observed that the rotamers of Asn140 in the Apo state may adopt a conformation favorable in the Apo state but not in the Holo state (see Results and Discussion and SI for more details). This could lead to sampling the Holo state with an incorrect orientation of Asn140, potentially introducing artifacts. The restriction (through a restraint) to one of these rotamers as an Apo endpoint must be accounted for. Therefore, a positional restraint was applied to the heavy atoms of the Asn140 residue and its neighboring atoms (all heavy atoms of the backbone and those within 6 Å of the ligand with a mild force constant of 2 ). This restraint was necessary to ensure that Asn140 and its neighboring atoms were properly sampled in both the Apo and Holo states, given the continuous switching of lambda between 0 (Apo) and 1 (Holo) in the Lambda-ABF-OPES simulations. Since both rotamers of Asn140 are equally favorable in the Apo state (see Fig. 2), an RTln(2) correction was added to the final computed . In addition, to avoid artifacts during alchemical decoupling as described in SI, the C atoms of BRD4 were restrained to their relaxed configuration.
@@
As all TI-based technique, Lambda-ABF requires the computation of potential derivatives with respect to the alchemical parameter which are not trivial to compute for many-body interactions such as polarization as present in the AMOEBA force field. In our previous work34 we used a simple interpolation of polarization between the end states which gives immediately the associated derivatives as the difference of these, but this is associated with an increase of the computational cost because of the need to solve 2 polarization equations at each timestep. Recently, analytical derivatives associated to a simple scaling of the polarizabilities and multipoles have been introduced49. Here we resorted to a more general formulation relying on the variational formulation of the many-body term and the Hellman-Feynman theorem: and
@@
because of the minimum conditions on the induced dipoles. Note that this formulation still holds for other many-body terms with similar variational formulation as is the case for fluctuating charges50, continuum solvation models51 or QM-MM52, 53. Further development of the gradients used in this work using Particle Mesh Ewald is given in Technical Appendix.
@@
In the following section, we will first provide a brief overview of the general binding modes of the 11 ligands in the BRD4 and Benzamidine-Trypsin system. We will then focus on the orientation of the Asn140 residue of BRD4, a key residue that directly interacts with the ligands, in both the Apo and Holo states, highlighting any significant differences. Finally, we will examine the binding affinities and correlate these findings with the experimental values.
@@
As depicted in Fig. 1(c), for BRD4 the 11 ligands are large, flexible, drug-like molecules, some of which are charged (ligand 1 and 4). These characteristics make them an ideal set for evaluating the performance of the new method. All ligands target a common binding site. Asparagine (Asn140), located within the BC-loop binding pocket (Fig. 1(a)), is the most critical residue, which interacts directly with the ligands. The side chain of Asn140 consists of an amide group () attached to a C, which is itself connected to the C of the backbone via rotatable bonds.
@@
In addition to Asn140, four conserved (polarizable) water molecules 54 in the binding site of BRD4 also play a crucial role in stabilizing the ligands within the binding pocket, thereby enhancing binding affinities. As illustrated in Fig. 1(a), one of these water molecules forms a bridge between the Tyr97 residue and the ligand (except for ligand 4). To preserve this stabilization, the X-ray water molecules were retained in the binding pocket during system preparation, as they contribute to a hydrogen-bond network involving the three other conserved water molecules and the protein.
The Benzamidine-Trypsin complex (see Fig. 1(d)) serves as another ideal case to evaluate the performance of the Lambda-ABF-OPES on a different target.
Using the equilibrated structures for both targets, we carried out absolute binding free energy calculations employing the novel Lambda-ABF-OPES method. The correlation plot between experimental values and calculated results of BRD4 is shown in Fig. 3(a) and reported in Table 1.
We observe a strong correlation between the experimental and calculated results, with a Pearson’s correlation coefficient (Pearson’s ) of 0.81. The analysis yielded a root mean square error (RMSE) of 1.1 kcal/mol and a mean absolute error (MAE) of 0.9 kcal/mol. These results are consistent with experimental data (see Fig. 3, panel (a)) and align well with those reported in Ref. 40 (see Fig. 3, panel (b)). For ligand 6, however, our obtained result aligned better with the experimental values. The key advantage of this method lies in its user-friendliness, reduced computational resource requirements, and rapid convergence. Compared to Lambda-ABF, the traditional fixed Lambda approach 41, and the newly developed CV-based approach 40, this method offers significantly improved computational efficiency while maintaining high accuracy. These attributes make it particularly well-suited for large-scale or high-throughput applications.
For the Benzamidine-Trypsin complex, we obtained an absolute binding free energy of kcal/mol. This result is in excellent agreement with the experimental value of kcal/mol 55, 56 and the findings of Ansari et al. 43. Moreover, achieving the same converged result in a significantly shorter simulation time further highlights the computational efficiency of the Lambda-ABF-OPES method.
@@
Figure 4 illustrates an example of a convergence plot over time for the ELE and VDW legs of the ligand 8-BRD4 complex phase using the Lambda-ABF and Lambda-ABF-OPES methods in a single replica (see SI Fig. S14-S18 for additional comparisons). In this case, the acceleration in convergence time is a factor of 9 for ELE and 4 for VDW. When considering different replicas, the convergence speedup for ELE ranges from 5 to 9 times, while for VDW, it ranges from 3 to 5 times.
@@
In this study, we introduced a new hybrid approach that combines Lambda-ABF with the exploratory version of the OPES method. This novel integration leverages the complementary strengths of ABF and OPESe to overcome critical limitations in alchemical free energy calculations, including inefficient exploration of configurational space and kinetic trapping in energy landscapes. By applying biases to the Lambda collective variable (CV) and incorporating the AMOEBA polarizable force field alongside DBC restraints, our method achieves unprecedented levels of sampling efficiency, up to nine times faster than the original Lambda-ABF technique.
Our application of this hybrid method to a diverse set of 11 drug-like molecules targeting BRD4 bromodomains and Benzamidine-Trypsin complex yielded close alignment between our computational results and experimental data. Importantly, this was achieved while using the highly accurate AMOEBA polarizable force field, demonstrating the feasibility of this approach for real-world drug discovery applications. This approach can be naturally extended to neural networks methodologies including Machine Learning Interatomic Potentials (MLIP) 64, 65 and Foundation models whose additional computational cost for free energy computations compared to FFs has limited to date their use in production.
@@
Data Availability Statement
The Tinker-HP and colvars input files used, and the structures are available on GitHub at: https://github.com/ansarinarjes/ABFE-L-ABF-OPES.git
@@
1 Supporting Information
The Supporting Information contains more detailed descriptions of the MD simulations and Lambda-ABF-OPES, including applied restraints, DBC selection, performance, and convergence analyses. In total, there are 3 tables and 23 figures with corresponding descriptions.